I'm really stuck on this particular question, which is:
A girl buys three apples and seven oranges for $4.56 .If an orange costs $0.28 more than an apple, how much does each fruit cost?
: I'm really stuck on this particular question, which is:
A girl buys three apples and seven oranges for $4.56 .If an orange costs $0.28 more than an apple, how much does each fruit cost?

Answer by Alan3354(1187)

I'm really stuck on this particular question, which is:
A girl buys three apples and seven oranges for $4.56 .If an orange costs $0.28 more than an apple, how much does each fruit cost?
0 solutions
-------------------
A = price of apples
R = price of oranges (O might look like a zero)
3A + 7R = 456
R = A + 28
-------
Substitute for R in eqn 1
3A + 7*(A+28) = 456
3A + 7A + 196 = 456
10A = 260
A = 26 cents
R = 26 + 28
R = 54 cents